OpenStudy (anonymous):

5 questions for online homework i need help!

OpenStudy (anonymous):

What is the fraction expressed in lowest terms?

OpenStudy (anonymous):

Okay what is the fraction? and what are the answer choices? @studydood66

OpenStudy (anonymous):

and there are no options i have to type the answer in

OpenStudy (anonymous):

@bibby

OpenStudy (anonymous):

i have a weird school /:

OpenStudy (anonymous):

Okay. So we have \[\frac{ 28 }{ 60 }\]

OpenStudy (anonymous):

Now can you tell that they are both even? @studydood66

OpenStudy (anonymous):

not realy

OpenStudy (anonymous):

oh yea i can

OpenStudy (anonymous):

Yes.. they are even! So we would divide 28 and 60 by 2... 28 divided by 2 is 14 and 60 divided by 2 is 30..

OpenStudy (anonymous):

So we have \[\frac{ 14 }{ 30 }\]

OpenStudy (anonymous):

BUT... We can simplify even more!

OpenStudy (anonymous):

but cant that go down further

OpenStudy (anonymous):

Lets divide by 2 again... 14 divided by 2 is 7 and 30 divided by 2 is 15 So the final answer is: \[\frac{ 7 }{ 15 }\]

OpenStudy (anonymous):

Great Job!

OpenStudy (anonymous):

Hope this helped! Have a great day! Also a medal would be much appreciated! Just click best response next to my answer. Thank You! @studydood66

OpenStudy (anonymous):

thats what i was typing

OpenStudy (anonymous):

i have 4 more

OpenStudy (anonymous):

Sorry... but you should probably post them in another post.

OpenStudy (anonymous):

So close this and put the others in another post.